OpenStudy (anonymous):

A triangle has sides measuring 3, 3√3, and 6 units. What are the angle measures of the triangle?

OpenStudy (anonymous):

@Darius25

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

and were is the triangle

OpenStudy (anonymous):

It doesn't show it. The answers are: A. 30°, 60°, 90° B. 45°, 45°, 90° C. 60°, 60°, 60° D. 30°, 52°, 60°

OpenStudy (anonymous):

This triangle is a half of a quadrilateral triangle, therefore, the angles measures are: 30, 60, 90.

OpenStudy (anonymous):

Thank you! Can you tell me how you figured it out?

OpenStudy (anonymous):

ill tell you

OpenStudy (anonymous):

he got it from the internet

OpenStudy (anonymous):

Okay?? Can you tell me how to do a similar problem?

OpenStudy (anonymous):

@thu1935

OpenStudy (anonymous):

There are many ways. a) use the right triangle identity: sise^2 + side^2 = hypotenuse^2 (3)^2 + (3V3)^2 = (6)^2 9 + 27 = 36 b) use triangle trigonomrtry -> sin A = 3/6 = 1/2 = sin 30 deg --> A = 30 deg sin B = 3V3/6 = V3/2 = sin 60 deg. --> B = 60 deg

OpenStudy (anonymous):

Thank you!